International Conference on Data Engineering (ICDE) Steering Committee Charter

January 1, 1996

1 Purpose of the ICDE Steering Committee

The objective of the International Conference on Data Engineering (ICDE) is to promote research and development of data engineering by: (a) providing a forum for the introduction and discussion of new research results, concepts, and technologies; (b) providing practicing engineers exposure to and an evaluation of evolving research, tools, and practices; (c) providing the research community exposure to the problems of practical applications of data engineering; and (d) encouraging the exchange of data engineering technologies and experience.

This conference attempts to meet the needs of a large and diverse constituency, which includes practitioners, researchers, educators, and users. The focus of the conference is on various aspects of data engineering that include, but are not limited to, the following areas: (a) data engineering tools and techniques; (b) parallel and distributed processing; (c) real-time distributed processing; (d) system architectures, integration, and modeling; (e) database design, modeling, and management;

(f) query design and implementation languages; (g) distributed database control;

(h) statistical databases; (i) algorithms for data and knowledge management; (j) performance evaluation of algorithms and systems; (k) data communications aspects; (l) system applications and experience; (m) knowledge-based systems; and (n) integrity, security, and fault tolerance.

The conference is managed by a Steering Committee, a Conference Organizing Committee, and a Program Committee. The functions of each committee are described in the following sections.

2 The Steering Committee

Purpose. To oversee the long-term planning and success of this conference, a Steering Committee plans for future conferences, evaluates how well each conference has achieved the objectives stated above, and implements improvements to better meet these objectives.

The purpose of the Steering Committee is to

1. provide a continuing committee to represent data engineering interests in the research community,

2. improve future ICDE meetings,

3. obtain feedback about ICDE for the purpose of improving the conference,

4. select the General and Program Chairs for future ICDE meetings,

5. select the sites and dates for future ICDE meetings, which may be in response to submitted proposals,

6. assist in the appointment of other chair positions in the Organizing and Program Committees.

The Steering Committee has the final authority on every aspect related to the organization and management of ICDE.

Composition of the Steering Committee. The Steering Committee is composed of

1. a chairperson,

2. Steering Committee members.

There is no limit on the number of Steering Committee members.

Chair of the Steering Committee. The Chair of the Steering Committee shall be elected from among the Steering Committee members for a term of the same duration as a Steering Committee member. The same Chair can be elected for one additional term. After finishing his/her term, the Chair must leave the Steering Committee for a minimum period of two years before becoming eligible to serve on the Steering Committee again.

Steering Committee Members. New members of the Steering Committee are selected by the current Steering Committee members in a Steering Committee meeting held at each ICDE. The committee is open to all participants of the conference who have significant experience in the operation of ICDE (such as serving as General and/or Program Chairs of past ICDEs). Any participant interested to serve should indicate his/her interest to the Steering Committee Chair before the meeting. The term of each member is two years, renewable once for another term of two years. After serving a four-year term, a Steering Committee member must be off the Steering Committee for a minimum period of two years before becoming eligible to serve again.

Steering Committee Meetings. A Steering Committee meeting must be held at each ICDE. If necessary, interim meetings can be called by the Steering Committee Chair through electronic mail or by other means.

Steering Committee members are expected to attend the Steering Committee meeting in each ICDE. Members who miss two consecutive meetings are considered withdrawn from the Committee.

Proxies. During meetings of the Steering Committee, each member shall have one vote. General proxy voting shall not be permitted, but on specific issues proxy voting shall be permitted if the Steering Committee Chair is informed beforehand of the intended vote by the absent member. Steering Committee members are to be informed sufficiently in advance of any known motions to be voted on during a forthcoming meeting.

3 Selection of Chair Positions

1. The General Chair of ICDE is selected by the Steering Committee and shall normally go to a new individual each time. This individual should be highly respected for his/her work in data engineering. In addition, he/she should have significant experience in organizing international conferences (e.g., as program committee members or program chairs of other conferences).

2. The name of the person recommended as the General Chair shall be part of the proposal (see Section 4). This person shall be from the country hosting the conference. If the conference is outside the United States, a General Co-Chair from the United States may also be appointed to serve as a liaison with the IEEE Computer Society. The Steering Committee shall be free to accept or reject any recommended names for General Chair in order to best promote the conference.

3. The Program Chair(s) are chosen by the General Chair with the advice and consent of the Steering Committee. It is especially important that the Program Chair(s) be active in the field, and have published frequently in data engineering journals and/or conferences. The Program Chair selects the program committee vice-chairs and members with the guidance of the General Chair.

4 Constituting a Conference

The Steering Committee obtains, through open and informal solicitations, suggestions for any or all of the following: possible conference locations, names of individuals who are suitable to serve as General Chair, and individuals who are suitable to serve as Program Chair(s). An individual or institution interested to propose and organize a future ICDE should follow the steps listed below.

1. Written proposals for conference locations should be sent to the Chair of the Steering Committee significantly before the proposed conference date. The location of the conference will generally be determined at least 18 months before the proposed date.

2. Proposals shall list facilities, venue, proposed management, financial arrangements, a brief summary of the conference structure, and any other information required by the Steering Committee. In particular, the scheduling of any other potentially conflicting data engineering-related meetings in that region within several months before or after the proposed meeting date must be evaluated as part of the proposal. Further, written commitments of financial support from sponsoring institutions, if any, should be submitted in the proposal. In general, the IEEE Computer Society will fully sponsor the conference, and only in exceptional situations that it will have a sponsorship of no less than 50%.

3. Those wishing to host a future ICDE should make their intentions known sufficiently in advance of the proposal date so that any information required by the Steering Committee can be obtained in a timely manner, and other activities can be integrated into a program of interest to all.

4. The Steering Committee seeks proposals for programs with broad appeal. Proposals can include tutorials, vendor displays, interesting and dynamic invited speakers, reports on workshops, and controversial panel discussions.

5. The Steering Committee reviews proposals by seeking advise from anonymous reviewers. Inputs from the IEEE Computer Society Technical Committee on Data Engineering will also be considered. If the reviews are positive, the Steering Committee will tentatively accept the proposal, and the proposer(s) will be asked to prepare a detailed budget using the IEEE Computer Society Technical Meeting Request Form (TMRF) and submit the form to the IEEE Computer Society and the Steering Committee Chair at least 15 months before the proposed conference date. Final approval of the TMRF will be made by the IEEE Computer Society.

5 General Chair and the Organizing Committee

In each conference, the General Chair, the Program Chair(s), and other selected chairs (for example, Tutorial Chair) form the Conference Organizing Committee for that conference and are charged with the planning, running, and oversight of that conference to ensure that it meets the objectives stated in Section 1.

The General Chair is responsible for the overall running of the conference, and serves as the Chair of the Organizing Committee. This Chair proposes the Organizing Committee plan, which must be approved by the Steering Committee.

The Organizing Committee approves all major decisions concerning the conference. The General Chair is responsible for the finances of the conference and for coordinating local arrangements. To these ends, the General Chair will appoint a Treasurer and a Local Arrangements Chair. The General Chair may also appoint a Publicity Chair, Publications Chair, Registration Chair and other positions as necessary.

The General Chair or the appointed Publicity Chair is responsible for publicity of the conference. All notices requesting papers, panels, tutorials and tool demonstrations as well as the advance program announcement and registration material shall be widely distributed in a timely manner. These include announcements in professional and research journals, the trade press, electronic bulletin boards and the World Wide Web, and separate mailings to individuals on (electronic and postal) mailing lists of interested or participating organizations. The goal is to allow full participation in all facets of the conference by the world-wide data engineering communities.

The General Chair is responsible for obtaining all necessary approvals and financial supports from conference sponsors.

The General Chair is responsible to solicit proposals on local facilities (such as hotels and universities) that host the conference. Hotel contracts must be ratified by the IEEE Computer Society, and the General Chair should not sign a hotel contract without first consulting the Conference Department of the IEEE Computer Society.

Special efforts should be made by the General and Program Committee chairs to get organizers of future ICDEs involved in the organizing and program committees of the present conference.

6 Program Chair(s) and the Program Committee

The Program Chair(s) shall be appointed no later than 18 months before the scheduled date of the conference. The Program Chair(s) are responsible for planning the technical program of the conference.

Program Committee. The Program Chair(s) will appoint individuals to serve in the program committee (with the advice and consent of the General Chair). The Program Committee should be diversified in terms of expertise, geographical distribution, race, and gender. Program Committee members shall be chosen on the basis of their technical qualifications and experience, and should include individuals with expertise in data engineering. The Program Chair(s) should promote continuity by keeping committee members from the previous conferences (e.g., members from past Organizing and Program Committees).

Program Committee Meetings. At least one-half of the program committee members should attend program committee meetings. In extraordinary situations and at the discretion of the Program Committee and General Chairs, these meetings may be conducted via telephone or electronic mail.

Call For Papers, etc. The Program Chair(s) will prepare the Call For Papers (with the Publicity Chair), send out notification of acceptance and rejection of submitted papers, and send out instructions to authors for submitting their final papers. The call for papers as well as individuals on the Program Committee shall encourage the submission of high quality papers that have not been previously published or currently submitted elsewhere, including reports on practical experiences with or evaluation of advanced technologies.

Reviews of Submitted Papers The Program Committee shall be responsible for reviewing all submitted papers, for submitting timely, informative reviews that provide authors with feedback about their papers, and for attending program committee meetings. The Program Chair(s) will determine the procedures for reviewing and accepting papers for the program. They shall distribute the papers for review, and arrange information so the papers can be discussed by knowledgeable reviewers during the program committee meeting. Program committee members can seek expert advice from external reviewers, but they shall be responsible for having reviewed the papers themselves. Each paper shall be reviewed by at least three reviewers and each paper shall be discussed at the program committee meeting.

The conference shall require full papers for review. Papers shall be reviewed by the program committee members. Direct discussion provides other committee members who might be familiar with the work an opportunity to contribute useful information that might affect the outcome of the review process. If consensus cannot be formed or there are additional concerns raised about a paper, additional information or reviews should be obtained.

Usually those papers that have received very poor reviews or very strong reviews can be discussed briefly, but an open discussion of all papers is required. It is important that the Program Committee consider all papers fairly.

If a committee member (Program, Organizing, or Steering Committee) of a forthcoming ICDE submits papers, care must be taken that his/her papers be reviewed at least as stringently as other submitted papers. Committee members shall leave the room during the discussion of their submitted papers or for any paper where he/she might have a conflict of interest.

All discussion about a paper shall remain in confidence. The names of reviewers are not to be revealed to authors. Reviewers, however, shall be free to identify themselves to an author if they so wish.

It is recognized that English may not be the native language of many of the authors of contributed papers. It is still expected, however, that all papers will be reasonably presented and written in acceptable English. It is strongly recommended that authors whose native language is not English to seek professional help in polishing the writing in their papers.

7 Amendments

Amendments to this Charter for the International Conference on Data Engineering require approval by two thirds of the members of the Steering Committee.

8 Revision History

The Bylaws of ICDE was modeled after the Bylaws of the IEEE-CS International Software Metrics Symposium and tailored by Benjamin Wah to the needs of ICDE in September 1995. The charter was ratified by the Steering Committee in November 1995.

Table 1: Current Steering Committee Members

Country  Name  Term  Period
Canada Nick Cercone  1 1/96 - 12/98
USA  Ahmed Elmagarmid 1 1/96 - 12/00
USA  Forouzan Golshani  1 1/96 - 12/99
Japan  Tadao Ichikawa  1 1/96 - 12/98
Germany Erich Neuhold  1 1/96 - 12/99
USA  Marek Rusinkiewicz 1 1/96 - 12/00
USA  Joseph Urban  1 1/96 - 12/97

Table 2: Retired Steering Committee Members

Country  Name  Period
USA  P. Bruce Berra (member)  1/84 - 12/89
USA  P. Bruce Berra (chair)  1/89 - 12/93
USA  John Carlis (member)  1/88 - 12/93
USA  Mike Liu (member)  1/89 - 12/93
USA  C. V. Ramamoorthy (chair) 1/84 - 12/88
USA Benjamin Wah (member) 1/87 - 12/93
USA  Benjamin W. Wah (Chair) 1/94 - 12/95